# Concept Steerers: Leveraging K-Sparse Autoencoders for Controllable Generations

### v1

@@ 3.3 Concept Steerers
Given a human-interpretable concept $C$ [^2] Defined by any user-provided prompt, e.g., “nudity”. we wish to steer, we first extract its text embedding $x_{C}$, pass it through k-SAE, and finally perform an element-wise addition with the input prompt embedding $x$. This can be expressed as:
@@
where $\lambda$ denotes a scalar that controls the steering strength. The steered vector $x_{\text{steered}}$ is used to condition the generation process. As we show in Sec. 4, our approach requires a k-SAE to be trained only once, and provides model-agnostic, fine-grained control over concept steering without degrading the overall generation quality.
@@ 4 Experiments
We first share the training setup followed by numerous results and ablations on concept steering.
Implementation details: We train k-sparse autoencoders on text embeddings with $k_{\text{aux}}=256$, and loss weight parameter $\alpha=1/32$ for 10k training steps. We train for a total training tokens of $400M$ on a batch size of $4096$ with the learning rate $0.0004$ using Adam (Kingma, 2014) optimizer. The k-SAE is trained with $k=32$ and an expansion factor of $4$, resulting in a total hidden size dimension $n=3072$ for Stable Diffusion (SD) 1.4 (Rombach et al., 2022) in the unsafe removal task. For style manipulation, we use $k=64$ with an expansion factor of $64$, resulting in a total hidden size dimension $n=49152$ for SD1.4 and $k=64$ with an expansion factor of $16$, resulting in a total hidden size dimension $n=32768$ for SDXL-Turbo (Sauer et al., 2025). These settings were found via ablation studies on downstream tasks and/or chosen based on overall training stability and sparsity. We apply a unit normalization constraint (Sharkey et al., 2023) on the decoder weights $W_{\text{dec}}$ of the k-SAE after each update. Although our method can be applied in an out-of-the-box manner to any text-to-image model, for a fair comparison with existing methods, we conduct experiments using SD 1.4 for unsafe concept removal and then expand our evaluation to more recent SDXL-Turbo and FLUX.1-dev. When steering concepts, we use a slightly modified encoder of k-SAE without the TopK activation function, as the TopK function clamps some possible important information that is critical for both maintaining visual quality and effective concept removal. More details in Appendix A.
@@ 4.1 Steering towards safety
Setup: First, we demonstrate the effectiveness of erasing unsafe concepts using k-SAEs. We use the Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023) to evaluate our method on steering nudity concepts and the Ring-A-Bell benchmark (Tsai et al., 2023) to assess steering performance on violent content. We train a k-SAE using features extracted from the residual stream of the 10th layer (out of 12 layers) of the text encoder in SD 1.4, on the prompts in I2P dataset, and evaluate its performance on both the I2P and Ring-A-Bell datasets. As in prior works (Tsai et al., 2023; Jain et al., 2024), we use nudity concept $C_{\text{nudity}}$ as “naked, nude, bare, exposed, stripped, topless, male genitalia, penis, buttocks” and use a slightly modified version for the violent concept $C_{\text{violence}}$ as “violence, blood”. We set steering strength $\lambda=-0.5$ for I2P dataset and $\lambda=-0.7$ for adversarial datasets including violent concept.
Evaluation metrics: To quantify the impact of our method on generation quality, we use FID (Heusel et al., 2017) and CLIP score (Hessel et al., 2021; Radford et al., 2021) on the COCO-30k dataset, evaluating 10k generated samples. We report Attack Success Rate (ASR), i.e., the percentage of generated images containing nudity or violence as a measure of how well a model reduces unsafe content generation. To this end, we use the NudeNet (Bedapudi, 2019) with a threshold of $0.45$ and Q16 violence detector (Schramowski et al., 2022), following prior work (Jain et al., 2024).
Baselines: We compare our method against inference-based approaches that do not require training or weight updates to the generative model, including SLD (Schramowski et al., 2023), SD with negative prompt (SD-NP), SAFREE (Yoon et al., 2024), and TraSCE (Jain et al., 2024). Additionally, we evaluate our method against training-based approaches, including ESD (Gandikota et al., 2023), FMN (Zhang et al., 2024), CA (Kumari et al., 2023), MACE (Lu et al., 2024), and SA (Heng & Soh, 2024), as well as approaches that require no training but involve weight updates, such as UCE (Gandikota et al., 2024) and RECE (Gong et al., 2025). We also try a variant of our model, where we steer in the opposite direction of the layer activation corresponding to the null text used for classifier-free guidance (Ho & Salimans, 2022), which we refer to as negative steering.
@@ 4.1.1 Steering nudity concept
Table 1: Performance comparison across different methods on I2P and COCO datasets. Lower ASR and FID indicate better performance; higher is better for CLIP. Our method achieves the lowest ASR by effectively removing nudity while preserving visual quality and prompt alignment. Bold: best. Underline: second-best. Gray : require training and weight updates, Pink : do not require training but update model weights, Blue : do not require either.
Method I2P COCO
ASR $\downarrow$ FID $\downarrow$ CLIP $\uparrow$
SDv1.4 17.80 16.71 31.3
ESD (Gandikota et al., 2023) 2.87 18.18 30.2
CA (Kumari et al., 2023) 1.04 24.12 30.1
MACE (Lu et al., 2024) 1.51 16.80 28.7
SA (Heng & Soh, 2024) 2.81 25.80 29.7
UCE (Gandikota et al., 2024) 0.87 17.99 30.2
RECE (Gong et al., 2025) 0.72 17.74 30.2
SLD-Max (Schramowski et al., 2023) 1.74 28.75 28.4
SLD-Strong (Schramowski et al., 2023) 2.28 24.40 29.1
SLD-Medium (Schramowski et al., 2023) 3.95 21.17 29.8
SD-NP 0.74 18.33 30.1
SAFREE (Yoon et al., 2024) 1.45 19.32 30.1
TraSCE (Jain et al., 2024) 0.45 17.41 29.9
Ours (w/o negative steering) 0.57 18.37 30.8
Ours 0.36 18.67 30.8
Figure 3: Qualitative comparisons of different approaches, including TraSCE and SAFREE, on the I2P dataset. Our method removes nudity without significantly altering the generated images, resulting in outputs that are better aligned with the input prompt.
Figure 4: Qualitative examples from the I2P dataset. Our method allows fine-grained control over the removal of specific concepts, removing only the intended concept while preserving the overall structure and style of the generated images.
@@
As shown in Table 1, our approach achieves state-of-the-art performance in steering unsafe concepts, yielding the lowest ASR ($\mathbf{0.36}$) on the I2P dataset and surpassing the previous best method. We note that incorporating negative steering slightly improves performance, demonstrating that our concept vector effectively models abstract concepts and, similar to negative prompting, yields a slight improvement in performance. Notably, our approach even outperforms both training-based methods (Gandikota et al., 2023; Kumari et al., 2023; Lu et al., 2024; Heng & Soh, 2024) and weight-updating methods (Gandikota et al., 2024; Gong et al., 2025), underscoring the effectiveness of our method. Furthermore, our method achieved the highest prompt-image correspondence, as indicated by the CLIP score on the COCO dataset ($30.8$), ranking just below the original SD 1.4 model ($31.3$). This is demonstrated in Fig. 3 and Fig. 4, where previous methods sometimes generate unrelated images when the prompt triggers unsafe content. By contrast, our method successfully removes nudity while preserving the overall structure and maintaining alignment with the input prompt. Moreover, as shown in Fig. 5, we demonstrate that our method can also steer the DiT-based (Peebles & Xie, 2023) FLUX (Labs, 2023) model in an out-of-the-box manner.
@@ 4.1.2 Steering violence concept
Table 2: Performance comparison across different methods on the Ring-A-Bell-Union (Violence) dataset. Lower values indicate better performance. Our method demonstrates competitive performance without compromising generation quality, as indicated by the FID scores in Table 1. Bold: best. Underline: second-best. Gray : require training and weight updates, Pink : do not require training but update model weights, Blue : do not require either.
Method Ring-A-Bell-Union
(Violence)$\downarrow$
SDv1.4 99.6
ESD (Gandikota et al., 2023) 86.0
FNM (Zhang et al., 2024) 98.8
CA (Kumari et al., 2023) 100.0
UCE (Gandikota et al., 2024) 89.8
RECE (Gong et al., 2025) 89.2
SLD-Max (Schramowski et al., 2023) 40.4
SLD-Strong (Schramowski et al., 2023) 80.4
SLD-Medium (Schramowski et al., 2023) 97.2
SD-NP 94.8
TraSCE (Jain et al., 2024) 72.4
Ours 43.7
We also evaluate our method’s performance in suppressing violent content generation, as presented in Table 2. As shown in Fig. 6, our method effectively reduces the generation of violent content compared to existing training-based and weight-update-based methods. Although SLD-Max achieves slightly better performance than ours, it significantly degrades overall image quality, yielding an FID of 28.75 compared to $\mathbf{18.67}$ for our approach (Table 1).
@@ 4.2 Steering of photographic styles and object attributes
Figure 7: Photographic style manipulation of SD 1.4 for the given prompt “geodesic landscape, john chamberlain, christopher balaskas, tadao ando, 4 k, ” where concept prompts are “minimalist” (Top) and “zoom-in, magnify” (Bottom), respectively. In the top row, the image is manipulated toward a maximalist style as $\lambda\rightarrow-1$, while it adopts a minimalist style as $\lambda\rightarrow 1$. Similarly, in the bottom row, the image appears zoomed out and becomes blurred as $\lambda\rightarrow-1$, whereas it becomes zoomed in and clearer as $\lambda\rightarrow 1$.
@@
Figure 8: Qualitative comparisons with weather Concept Sliders on SDXL-Turbo. Note that Concept Sliders train specific sliders: winter weather slider and a dark weather slider, whereas our method trains a k-SAE only once for different concepts. Top: “A photo of a tree with a bench, realistic, 8k” with concept to steer = “winter.” Bottom: “A photo of a forest, realistic, 8k” with the concept to steer = “low light.” Notice how in the top image our method also removes leaves while in the bottom image, our method effectively applies a low-light effect to the original image.
As shown in Fig. 7 and Fig. 8 we can adjust its photographic style, including “zoom-in” and “minimalist.” In Fig. 8, we compare our results with Concept Sliders (Gandikota et al., 2025) on SDXL-Turbo where Concept Sliders train separate models for each weather condition style. Remarkably, our method can effectively steer concepts like weather conditions and photographic styles. We note that I2P dataset in addition to the semantic concepts such as nudity and violence, also had descriptors about general photographic styles such as “full shot” or seasons “winter”. We believe that k-SAE internalized these concepts offering us a powerful tool to surgically steer them. This powerful result highlights the generalizable capability of k-SAEs to learn diverse monosemantic concepts. This is corroborated by our results in Fig. 9, where we show that our method can manipulate image compositions, changing a close-up image of a dog into a “full shot” of a dog while preserving the appearance of its head part.
@@
Figure 9: Image composition manipulation using SDXL-Turbo for the prompt “A dog” with the concept prompt “Full shot.” Notice how as $\lambda\rightarrow 1$, the generated image transitions from a close-up of the face to a full shot.
Figure 10: Object attribute manipulation of SDXL-Turbo for the given prompts “A car” (Top) and “A photo of a tree” (Bottom), where the concept prompts are “A blue car” (Top) and “Tree with autumn leaves” (Bottom). By adjusting $\lambda$, our method transitions the image toward the desired concept specified by the prompts.
@@
Finally, in Fig. 10, we use the same k-SAE to effectively manipulates object attributes. Here, we inject a concept for an object present in the image, such as “blue [object]” or “tree with autumn leaves.” We note that the resulting generations preserve most of the original content while successfully injecting the desired concept. These results demonstrate the universal applicability of a k-SAE without the need to train separate adapters for each concept. We wish to continue exploring the limits of universality of k-SAEs in the future.

### v3

@@ 3.3 Concept Steerers
Given a human-interpretable concept $C$ we wish to steer, we first extract its text embedding $x_{C}$, pass it through the encoder and decoder of k-SAE, and finally perform an element-wise addition of the reconstructed k-SAE output with the input prompt embedding xx. [^2] CC is defined by any user-provided prompt, e.g., “nudity.” We express this as:
@@
where $\lambda$ denotes a scalar that controls the steering strength. The steered vector $x_{\text{steered}}$ is used to condition the generation process during inference. As we show in Sec. 4, our approach requires a k-SAE to be trained only once, and provides model-agnostic, fine-grained control over concept steering without degrading the overall generation quality.
@@ 4 Experiments
Implementation details: We train k-sparse autoencoders on text embeddings with $k_{\text{aux}}=256$, and loss weight parameter $\alpha=1/32$ for 10k training steps. We train for a total training tokens of $400M$ on a batch size of $4096$ with the learning rate $0.0004$ using Adam kingma2014adam optimizer. The k-SAE is trained with $k=32$ and an expansion factor of $4$, resulting in a total hidden size dimension $n=3072$ for Stable Diffusion (SD) 1.4 rombach2022high in the unsafe removal task. For style manipulation, we use $k=64$ with an expansion factor of $64$, resulting in a total hidden size dimension $n=49152$ for SD 1.4 and $k=64$ with an expansion factor of $16$, resulting in a total hidden size dimension $n=32768$ for SDXL-Turbo sauer2025adversarial . These settings were found via ablation studies on downstream tasks and/or chosen based on overall training stability and sparsity. We apply a unit normalization constraint sharkey2023taking on the decoder weights $W_{\text{dec}}$ of the k-SAE after each update. Although our method can be applied in a model-agnostic manner to any text-to-image model, for a fair comparison with existing methods, we conduct experiments using SD 1.4 for unsafe concept removal and then expand our evaluation to more recent SDXL-Turbo and FLUX.1-dev flux2023 . More details in Appendix A.
Motivation to steer only text-embeddings: We note that we train the k-SAE only once on text embeddings. Our choice was motivated by several reasons: (a) steering only via text embeddings offers more control to end users by having them simply specify the concept they wish to steer via text prompts, (b) most text-to-image model architectures vary in the visual training data and the diffusion model architecture but primarily use a handful of text encoders radford2021learning ; cherti2023reproducible ; raffel2020exploring . Thus, Concept Steerer offers a way to reuse a k-SAE across different generative models that use the same text encoder.
@@ 4.1 Steering towards safety
Figure 3: (a) Qualitative comparison of different approaches, including SAFREE yoon2024safree and TraSCE jain2024trasce , on the I2P dataset. Our method removes nudity without significantly altering the generated images, resulting in outputs better aligned with the input prompt. (b) Qualitative examples from the I2P dataset. Our method allows fine-grained control over the removal of specific concepts, removing only the intended concept while preserving the overall structure and style of the generated images. (c) Qualitative examples from the Ring-A-Bell dataset. Our method successfully removes the abstract concept of violence, as shown by the absence of blood in the right images. The images are intentionally blurred for display purposes as they are disturbing.
Setup: First, we demonstrate the effectiveness of erasing unsafe concepts using k-SAEs. We use the Inappropriate Image Prompts (I2P) dataset schramowski2023safe to evaluate our method on steering nudity concepts and the Ring-A-Bell benchmark tsai2023ring to assess steering performance on violent content. For comparing different techniques, each image is generated using the exact seed and the CFG scale specified in both datasets. We train a k-SAE using features extracted from the residual stream of the 10th layer (out of 12 layers) of the text encoder in SD 1.4, on the prompts in I2P dataset, and evaluate its performance on both the I2P and Ring-A-Bell datasets. As in prior works jain2024trasce ; tsai2023ring , we use nudity concept $C_{\text{nudity}}$ as “naked, nude, bare, exposed, stripped, topless, male genitalia, penis, buttocks” and use a slightly modified version for the violent concept $C_{\text{violence}}$ as “violence, blood”. We set steering strength $\lambda=-0.5$ for I2P dataset and $\lambda=-0.7$ for adversarial datasets including violent concept.
Evaluation metrics: To quantify the impact of our method on generation quality, we use FID heusel2017gans and CLIP score hessel2021clipscore ; radford2021learning on the COCO-30k dataset, evaluating 10k generated samples. We report Attack Success Rate (ASR), i.e., the percentage of generated images containing nudity or violence as a measure of how well a model reduces unsafe content generation. To this end, we use the NudeNet bedapudi2019nudenet with a threshold of $0.45$ and Q16 violence detector schramowski2022can , following prior work jain2024trasce .
Baselines: We compare our method against inference-based approaches that do not require training or weight updates to the generative model, including SDID li2024self , SLD schramowski2023safe , SD with negative prompt (SD-NP), SAFREE yoon2024safree , and TraSCE jain2024trasce . Additionally, we evaluate our method against training-based approaches, including ESD gandikota2023erasing , FMN zhang2024forget , CA kumari2023ablating , MACE lu2024mace , and SA heng2024selective , as well as approaches that require no training but involve weight updates, such as UCE gandikota2024unified and RECE gong2025reliable . We follow publicly available implementations exactly. More implementation details are in Appendix A.
@@ 4.1.1 Steering nudity concept
Table 1: Attack Success Rate (ASR) of different methods on I2P and various adversarial attack datasets. Lower ASR and FID indicate better performance; higher is better for CLIP. Our method achieves the lowest ASR on I2P and best overall robustness on average (column title: AVG) across 44 datasets (Ring-a-Bell, MMA-Diffusion, P4D, UnlearnDiffAtk), by effectively removing nudity implicitly embedded in the model. We also show that our model competes well on preserving visual quality (lower FID score) and prompt alignment (higher CLIP score). Bold: best. Underline: second-best. Gray : require training and weight updates, Pink : do not require training but update model weights, Blue : do not require either. Best viewed in color.*Due to the small size of both datasets, the performance gap relative to other models amounts to only a few images (1–7).
Method I2P $\downarrow$ Ring-A-Bell $\downarrow$ MMA-Diffusion $\downarrow$ P4D $\downarrow$ UnLearnDiffAtk $\downarrow$ Avg $\downarrow$ COCO
K77 K38 K16 Avg FID $\downarrow$ CLIP $\uparrow$
SD 1.4 (rombach2022high, ) 17.8 85.26 87.37 93.68 88.10 95.70 98.70 69.70 87.05 16.71 31.3
ESD (gandikota2023erasing, ) 2.87 20.00 29.47 35.79 28.42 12.70 9.27 15.49 16.47 18.18 30.2
SA (heng2024selective, ) 2.81 63.15 56.84 56.84 58.94 9.30 47.68 12.68 32.15 25.80 29.7
CA (kumari2023ablating, ) 1.04 86.32 91.69 94.26 90.76 9.90 10.60 5.63 29.22 24.12 30.1
MACE (lu2024mace, ) 1.51 2.10 0.00 0.00 0.70 0.50 2.72 2.82 1.68 16.80 28.7
UCE (gandikota2024unified, ) 0.87 10.52 9.47 12.61 10.87 17.70 29.93 9.86 17.09 17.99 30.2
RECE (gong2025reliable, ) 0.72 5.26 4.21 5.26 4.91 13.30 21.77 5.63 11.40 17.74 30.2
SDID (li2024self, ) 3.77 94.74 95.79 90.53 93.68 34.20 69.54 30.99 57.10 22.16 31.1
SLD-Max (schramowski2023safe, ) 1.74 23.16 32.63 42.11 32.63 52.10 35.76 9.15 32.41 28.75 28.4
SLD-Strong (schramowski2023safe, ) 2.28 56.84 64.21 61.05 60.70 61.30 49.01 18.31 47.33 24.40 29.1
SLD-Medium (schramowski2023safe, ) 3.95 92.63 88.42 91.05 90.70 65.70 68.21 33.10 64.42 21.17 29.8
SD-NP 0.74 17.89 40.42 34.74 31.68 44.44 24.00 7.80 26.98 18.33 30.1
SAFREE (yoon2024safree, ) 1.45 35.78 47.36 55.78 46.31 36.40 40.82 10.56 33.52 19.32 30.1
TraSCE (jain2024trasce, ) 0.45 1.05 2.10 2.10 1.75 16.60 3.97 0.70 5.75 17.41 29.9
Ours 0.36 3.16* 8.42* 9.47* 7.02* 6.00 1.99 2.11* 4.28 18.67 30.8
Figure 4: Qualitative examples from the I2P dataset with FLUX. Our method is model-agnostic and can be applied to both U-Net-based SD 1.4 and SDXL-Turbo, as well as DiT-based FLUX.
@@
As shown in Table 1, our approach achieves state-of-the-art performance in steering unsafe concepts, yielding the lowest ASR ($\mathbf{0.36}$) on the I2P dataset and surpassing the previous best method. Notably, our approach even outperforms both training-based methods gandikota2023erasing ; heng2024selective ; kumari2023ablating ; lu2024mace and weight-updating methods gandikota2024unified ; gong2025reliable , underscoring the effectiveness of our method. Furthermore, our method achieved one of the highest prompt-image correspondences, as indicated by the CLIP score on the COCO dataset ($30.8$), closely following the original SD 1.4 model ($31.3$). This is demonstrated in Fig. 3 (a) and (b), where previous methods sometimes generate unrelated images when the prompt triggers unsafe content. By contrast, our method successfully removes nudity while preserving the overall structure and maintains alignment with the input prompt. Moreover, as shown in Fig. 4, we demonstrate that our method is not limited to a single generative architecture and can also steer the DiT-based peebles2023scalable FLUX flux2023 model in a model-agnostic manner.
@@ 4.1.2 Steering violence concept
Table 2: Attack Success Rate (ASR) on different methods on the Ring-A-Bell-Union (Violence) dataset. Lower values indicate better performance. Our method demonstrates competitive performance without compromising generation quality, as indicated by the FID scores in Table 1. Bold: best. Underline: second-best. Gray : require training and weight updates, Pink : do not require training but update model weights, Blue : do not require either. Best viewed in color.
SD 1.4 rombach2022high ESD gandikota2023erasing FNM zhang2024forget CA kumari2023ablating UCE gandikota2024unified RECE gong2025reliable SLD-Max schramowski2023safe SLD-Strong schramowski2023safe SLD-Medium schramowski2023safe SD-NP TraSCE jain2024trasce Ours
Ring-A-Bell-Union (Violence) $\downarrow$ 99.6 86.0 98.8 100.0 89.8 89.2 40.4 80.4 97.2 94.8 72.4 43.7
We also evaluate our method’s performance in suppressing violent content generation, as presented in Table 2. Note that FID heusel2017gans and CLIP score hessel2021clipscore are computed on COCO dataset lin2014microsoft for a given model, so the scores reported in Table 1 hold true for the same models in Table 2. As shown in Fig. 3 (c), our method effectively reduces the generation of violent content compared to existing training-based and weight-update-based methods. Although SLD-Max schramowski2023safe achieves slightly better performance than ours, it significantly degrades overall image quality, yielding an FID of $28.75$ compared to $\mathbf{18.67}$ for our approach (Table 1, more visual examples in Appendix Fig. 11).
@@ 4.2 Steering of photographic styles and object attributes
Figure 5: Photographic style manipulation of SD 1.4 for the given prompt “geodesic landscape, john chamberlain, christopher balaskas, tadao ando, 4 k,” where concept prompts are “minimalist” (Left) and “zoom-in, magnify” (Right), respectively. On the left, the image is manipulated towards a maximalist style as $\lambda\rightarrow-1$, while it adopts a minimalist style as $\lambda\rightarrow 1$. Similarly, on the right, the image appears zoomed out and becomes blurred as $\lambda\rightarrow-1$, whereas it becomes zoomed in and clearer as $\lambda\rightarrow 1$.
@@
Figure 6: Qualitative comparisons with weather Concept Sliders on SDXL-Turbo. Concept Sliders train specific sliders (e.g., winter, dark), whereas ours trains a k-SAE once for multiple concepts. Left: “tree with bench” with “winter” concept. Right: “forest” with “low light.” Our method removes leaves and applies a low-light effect.
@@
Figure 7: Concept manipulation using SDXL-Turbo. (a) Image composition manipulation for the prompt “A dog” with the concept prompt “Full shot.” As $\lambda\rightarrow 1$, the image shifts from a close-up to a full-body view. (b) Object attribute manipulation for the prompts “A car,” where the concept prompts are “A blue car.” By adjusting $\lambda$, our method transitions the image toward the desired concept.
As shown in Fig. 5 we can adjust its photographic style, including “zoom-in” and “minimalist.” In Fig. 6, we compare our results with Concept Sliders gandikota2025concept on SDXL-Turbo where Concept Sliders train separate models for each weather condition style. Remarkably, our method can effectively steer concepts like weather conditions and photographic styles. We note that I2P dataset in addition to the semantic concepts such as nudity and violence, also had descriptors about general photographic styles such as “full shot” or seasons “winter.” We believe that k-SAE internalized these concepts offering us a powerful tool to surgically steer them. This powerful result highlights the generalizable capability of k-SAEs to learn diverse monosemantic concepts. This is corroborated by our results in Fig. 7 (a), where we show that our method can manipulate image compositions, changing a close-up image of a dog into a “full shot” of a dog while preserving the appearance of its head part.
@@
Finally, we use the same k-SAE to effectively manipulate object attributes. Here, we inject a concept for an object present in the image, such as “blue [object].” We note that the resulting generations preserve most of the original content while successfully injecting the desired concept (Fig. 7 (b)). These results demonstrate the universal applicability of a k-SAE without the need to train separate adapters gandikota2025concept or training data li2024self for each concept to control.